# Equipping Sketch Patches with Context-Aware Positional Encoding for Graphic Sketch Representation

### v1

@@ 3.3 Training a DC-gra2seq
Table 1: Controllable sketch synthesis performance (%). “@$k$” indicates the top-$k$ accuracy.
Model DS1 DS2 DS3
$Rec\uparrow$ $Ret\uparrow$ $Rec\uparrow$ $Ret\uparrow$ $Rec\uparrow$ $Ret\uparrow$
@1 @10 @50 @1 @10 @50 @1 @10 @50
sketch-pix2seq [11] 83.99 13.45 30.12 49.99 85.46 50.94 71.38 80.15 79.13 22.92 47.55 58.19
Song et al. [3] 91.77 16.41 36.43 52.22 86.98 58.84 76.84 80.06 83.28 25.47 43.39 56.16
Xu et al. [5] 93.32 22.70 54.40 75.04 90.01 51.98 75.79 83.26 82.34 25.25 54.04 72.90
RPCL-pix2seq [12] 93.18 17.86 38.87 55.30 88.73 53.19 71.60 87.91 81.80 28.80 59.05 77.52
RPCL-pix2seqH [13] 94.97 71.36 90.29 94.86 92.62 85.91 93.67 95.90 87.82 81.22 92.15 94.90
SketchHealer [1] 91.04 58.80 82.15 91.33 94.04 87.54 96.19 98.26 87.03 68.52 82.37 86.57
SketchHealer 2.0 [2] 93.13 57.19 84.54 90.26 90.94 87.37 94.59 97.60 87.37 50.67 76.11 82.42
SketchLattice [19] 75.91 6.55 14.01 26.72 71.80 6.91 14.76 28.82 62.21 5.90 10.36 19.39
SketchLattice+ 95.18 72.74 91.60 97.14 94.30 90.56 97.78 99.27 89.49 87.27 96.82 98.98
SP-gra2seq [8] 95.91 94.88 99.11 99.72 94.85 90.83 98.29 99.08 89.83 94.05 98.72 99.57
DC-gra2seq 96.26 96.84 99.48 99.88 95.70 92.75 98.08 99.42 90.41 96.27 99.47 99.81
@@ 4.1 Preparations
Datasets. Three datasets from QuickDraw [15] are selected, and all of them are adopted from [8] to make a direct comparison with the state-of-the-art SP-gra2seq [8]. Dataset 1 (DS1), which contains bee, bus, flower, giraffe and pig, evaluates representing performance on large variations of sketches within the same category. Dataset 2 (DS2), which collects airplane, angel, apple, butterfly, bus, cake, fish, spider, the Great Wall and umbrella, evaluates whether the models are sensitive to categorical patterns. Dataset 3 (DS3) collects the five categories in DS1 along with three new ones (car, cat and horse). Sketches in different categories are with shared stylistic patterns, e.g., both giraffes and horses can head left or right. Each category contains $70K$ training, $2.5K$ valid and $2.5K$ test sketches ($10001K=$ 1000).
Baselines. We compare DC-gra2seq with ten baseline models. Song et al. [3] and Xu et al. [5] learn representations from sketch images and their corresponding sketch sequences simultaneously. Sketch-pix2seq [11] learns representations from sketch images only. We make them a group to verify the advantage brought by drawing orders as an additional source. RPCL-pix2seq [12] and RPCL-pix2seqH [13] both constrain sketch codes with a specific distribution. SketchLattice [19], SP-gra2seq [8], SketchHealer [1] and SketchHealer 2.0 [2] learn graphic sketch representations. The sequential information from drawing orders are not used in SketchLattice and SP-gra2seq, while they are treated as prior knowledge to construct graph edges in SketchHealer and SketchHealer 2.0. Moreover, following [8], we provide SketchLattice+ to replace the original node embeddings captured from coordinates on lattice with the ones from sketch patches, to make a fair comparison.
@@ 4.2 Controllable sketch synthesis
Controllable sketch synthesis requires a model to generate sketches with the expected categorical and stylistic patterns. Table 1 reports the quantitative performances on three datasets.
@@
Table 2: Sketch healing performance (%). “Mask” denotes the probability for creating masks. “SH”, “SL”, “SP” and “DC” indicate SketchHealer, SketchLattice, SP-gra2seq and the proposed DC-gra2seq, respectively.
Mask Model DS1 DS2 DS3
$Rec\uparrow$ $Ret\uparrow$ $Rec\uparrow$ $Ret\uparrow$ $Rec\uparrow$ $Ret\uparrow$
@1 @10 @50 @1 @10 @50 @1 @10 @50
10% SH [1] 70.38 14.25 27.91 45.51 70.56 35.22 55.86 66.24 60.91 15.10 37.89 53.10
SH 2.0 [2] 90.18 19.45 41.51 59.94 87.07 39.40 64.01 81.20 76.45 15.55 39.99 61.67
SL [19] 54.18 1.09 5.48 14.00 52.97 1.34 6.86 17.41 44.14 0.71 4.19 11.36
SL+ 89.98 19.86 43.59 63.70 90.52 46.98 70.21 83.12 78.28 22.56 44.89 62.11
SP [8] 92.90 41.24 65.74 80.16 91.24 50.42 73.35 85.18 83.38 40.20 63.40 77.52
DC (ours) 94.45 74.59 88.83 94.31 92.64 67.61 84.31 91.17 85.21 71.66 86.34 92.35
@@
SH 2.0 [2] 79.05 4.66 14.54 28.44 75.08 10.05 24.51 39.57 60.74 3.75 11.66 22.36
SL [19] 32.03 0.08 2.05 5.96 31.73 0.88 3.64 9.25 23.06 0.41 2.41 5.96
SL+ 70.91 1.02 5.26 14.28 81.76 10.43 26.44 42.90 67.31 2.75 11.13 23.68
SP [8] 84.85 12.87 29.39 45.58 82.85 12.19 29.37 46.53 71.07 5.65 17.40 32.90
DC (ours) 86.50 30.30 51.38 65.84 84.62 29.92 49.26 62.44 72.84 27.64 47.34 60.61
Figure 3: Qualitative comparisons on sketch healing.
@@ 4.3 Sketch healing
Table 2 presents the sketch healing results on three datasets. Our DC-gra2seq yields significant improvements on $Ret$. When dealing with sketch patches corrupted by masks, the passing messages from the semantic-related patches by self-attention scores could contain valuable missing information under the masks. Moreover, by taking account of PEs, DC-gra2seq might realize where these semantic-related patches are drawn from the contextual positions in a drawing order. By incorporating both visual patterns from the canvas and their sequential information from the embedded PEs, DC-gra2seq learns much more comprehensive and accurate sketch representations.
@@ 4.4 Variant-drawing-protected DC-gra2seq
Table 4: Controllable sketch synthesis (Mask = 0%) and sketch healing (Mask = 10%, 30%) performances of DC-gra2seq by allowing PEs to participate in the construction of graph nodes, edges or not. “Node $\checkmark$ Edge $\times$” denotes the proposed DC-gra2seq.
Mask Node Edge DS1 DS2 DS3
$Rec$ $Ret$ $Rec$ $Ret$ $Rec$ $Ret$
@1 @10 @50 @1 @10 @50 @1 @10 @50
0% $\times$ $\times$ 94.51 80.90 95.09 98.19 85.84 41.00 71.80 87.31 83.93 57.50 83.52 93.45
$\checkmark$ $\times$ 96.26 96.84 99.48 99.88 95.70 92.75 98.08 99.42 90.41 96.27 99.47 99.81
$\times$ $\checkmark$ 95.69 79.22 93.92 97.77 92.41 67.49 88.30 95.30 88.48 76.25 92.64 97.24
$\checkmark$ $\checkmark$ 95.91 95.66 99.23 99.78 94.67 90.00 97.88 99.32 89.05 94.74 99.44 99.75
@@
$\checkmark$ $\times$ 94.45 74.59 88.83 94.31 92.64 67.61 84.31 91.17 85.21 71.66 86.34 92.35
$\times$ $\checkmark$ 93.48 50.64 75.10 87.46 89.48 45.24 70.13 83.33 84.79 46.77 70.37 82.90
$\checkmark$ $\checkmark$ 93.96 70.56 86.10 92.56 91.55 67.10 84.08 91.02 85.18 69.96 86.67 92.43
@@
$\checkmark$ $\times$ 86.50 30.30 51.38 65.84 84.62 29.92 49.26 62.44 72.84 27.64 47.34 60.61
$\times$ $\checkmark$ 84.23 16.28 36.80 55.01 80.40 18.62 37.78 53.70 71.99 14.56 31.54 47.96
$\checkmark$ $\checkmark$ 85.00 26.15 45.94 61.06 83.94 29.38 48.55 62.22 72.26 25.88 46.33 59.82
@@ 4.5 The impact of injecting drawing orders into nodes or edges
We further discuss the impact of injecting sketch drawing orders into graph nodes, edges or both via PEs. To make PEs access to edge construction, we adopt the computation of self-attention scores in vanilla self-attention network [10], i.e., $\alpha^{\text{atten}}(i,j)=(\bm{v}_{ti}+\bm{P}(i))^{\top}(\bm{v}_{tj}+\bm{P}(j))$. A softmax activation function is applied over $\bm{A}_{t}$ in Eq. (6) before sent into Eq. (8) for message aggregation.
Table 4 presents the experimental results on three datasets. When PEs attend the edge construction only (Node $\times$, Edge $\checkmark$), it outperforms the version which never uses PEs (Node $\times$, Edge $\times$) in most cases. The usage of drawing orders via PEs embedded graph edges is beneficial to improve sketch learning, but the performance gain is sensitive to various sketch drawings, e.g., it significantly outperforms “Node $\times$, Edge $\times$” on $Rec$ in DS2, but is defeated on top-1, top-10 $Ret$ in DS1. Similarly, our DC-gra2seq (Node $\checkmark$, Edge $\times$) outperforms “Node $\checkmark$, Edge $\checkmark$” in most cases. As a result, embedding drawing orders into graph nodes is a proper way to make better use of sequential information.

### v2

@@ 4.1 Preparations
Datasets. Three datasets from QuickDraw [16] are selected, and all of them are adopted from [7] to make a direct comparison with SP-gra2seq [7]. Dataset 1 (DS1), which contains bee, bus, flower, giraffe and pig, evaluates representing performance on large variations of sketches within the same category. Dataset 2 (DS2), which collects airplane, angel, apple, butterfly, bus, cake, fish, spider, the Great Wall and umbrella, evaluates whether the models are sensitive to categorical patterns. Dataset 3 (DS3) collects the five categories in DS1 along with three new ones (car, cat and horse). Sketches in different categories are with shared stylistic patterns, e.g., both giraffes and horses can head left or right. Each category contains $70K$ training, $2.5K$ valid and $2.5K$ test sketches ($10001K=$ 1000).
Baselines. We compare DC-gra2seq with eleven baseline models. Song et al. [3] and Xu et al. [5] learn representations from sketch images and their corresponding sketch sequences simultaneously. Sketch-pix2seq [11] and Lmser-pix2seq [14] learn representations from sketch images. We make them a group to verify the advantage brought by drawing orders as an additional source. RPCL-pix2seq [12] and RPCL-pix2seqH [13] both constrain sketch codes with a specific distribution. SketchLattice [20], SP-gra2seq [7], SketchHealer [1] and SketchHealer 2.0 [2] learn graphic sketch representations. The sequential information from drawing orders are not used in SketchLattice and SP-gra2seq, while they are treated as prior knowledge to construct graph edges in SketchHealer and SketchHealer 2.0. Moreover, following [7], we provide SketchLattice+ to replace the original node embeddings captured from coordinates on lattice with the ones from sketch patches, to make a fair comparison.
@@ 4.2 Controllable sketch synthesis
Controllable sketch synthesis requires a model to generate sketches with the expected categorical and stylistic patterns. Table 1-3 report the quantitative performances on three datasets, respectively.
Table 1: Controllable sketch synthesis performance (%) on DS1. “@$k$” indicates the top-$k$ accuracy.
Model $Rec\uparrow$ $Ret\uparrow$
@1 @10 @50
sketch-pix2seq [11] 83.99 13.45 30.12 49.99
Song et al. [3] 91.77 16.41 36.43 52.22
Xu et al. [5] 93.32 22.70 54.40 75.04
RPCL-pix2seq [12] 93.18 17.86 38.87 55.30
RPCL-pix2seqH [13] 94.97 71.36 90.29 94.86
SketchHealer [1] 91.04 58.80 82.15 91.33
SketchHealer 2.0 [2] 93.13 57.19 84.54 90.26
SketchLattice [20] 75.91 6.55 14.01 26.72
SketchLattice+ 95.18 72.74 91.60 97.14
Lmser-pix2seq [14] 95.26 92.90 97.72 99.02
SP-gra2seq [7] 95.91 94.88 99.11 99.72
DC-gra2seq 96.26 96.84 99.48 99.88
Table 2: Controllable sketch synthesis performance (%) on DS2.
Model $Rec\uparrow$ $Ret\uparrow$
@1 @10 @50
sketch-pix2seq [11] 85.46 50.94 71.38 80.15
Song et al. [3] 86.98 58.84 76.84 80.06
Xu et al. [5] 90.01 51.98 75.79 83.26
RPCL-pix2seq [12] 88.73 53.19 71.60 87.91
RPCL-pix2seqH [13] 92.62 85.91 93.67 95.90
SketchHealer [1] 94.04 87.54 96.19 98.26
SketchHealer 2.0 [2] 90.94 87.37 94.59 97.60
SketchLattice [20] 71.80 6.91 14.76 28.82
SketchLattice+ 94.30 90.56 97.78 99.27
Lmser-pix2seq [14] 94.27 89.95 97.08 99.25
SP-gra2seq [7] 94.85 90.83 98.29 99.08
DC-gra2seq 95.70 92.75 98.08 99.42
Table 3: Controllable sketch synthesis performance (%) on DS3.
Model $Rec\uparrow$ $Ret\uparrow$
@1 @10 @50
sketch-pix2seq [11] 79.13 22.92 47.55 58.19
Song et al. [3] 83.28 25.47 43.39 56.16
Xu et al. [5] 82.34 25.25 54.04 72.90
RPCL-pix2seq [12] 81.80 28.80 59.05 77.52
RPCL-pix2seqH [13] 87.82 81.22 92.15 94.90
SketchHealer [1] 87.03 68.52 82.37 86.57
SketchHealer 2.0 [2] 87.37 50.67 76.11 82.42
SketchLattice [20] 62.21 5.90 10.36 19.39
SketchLattice+ 89.49 87.27 96.82 98.98
Lmser-pix2seq [14] 89.03 89.21 97.63 99.12
SP-gra2seq [7] 89.83 94.05 98.72 99.57
DC-gra2seq 90.41 96.27 99.47 99.81
@@
Moreover, Fig. 3 also presents the qualitative comparisons on controllable sketch synthesis. SketchHealer, SketchHealer 2.0, SketchLattice+, Lmser-pix2seq and SP-gra2seq are selected as baselines, as they achieve high performances in quantitative comparisons. The sketches generated by DC-gra2seq could preserve more detailed patterns from inputs as the ground truth, e.g., the shape of the bus (in 2-nd column) and the candles on the cake (in 8-th column).
Figure 3: Qualitative comparisons on controllable sketch synthesis.
We also reveal the latent space learned by DC-gra2seq on DS3, shown in Fig. 4(a). Each dot indicates a captured latent code of a sketch and its color denotes the corresponding category. The high-dimensional codes are mapped into a two-dimensional space via t-SNE [36]. We can figure out that the sketches in the same category are clustered in a specific latent region, making categorical patterns easier to be controlled in sketch generating process. Besides, Fig. 4(b) presents sketches generated by interpolated latent codes. More specifically, in each row, the latent codes are computed via a linear interpolation between the ones at two ends. The characteristics of sketches are gradually changed along with the movement of latent codes from one end to the other. It indicates that sketch patterns are positioned smoothly in latent space, which enables DC-gra2seq to flexibly control sketch patterns.
Figure 4: Generating sketches by DC-gra2seq with interpolated latent codes. (a) The latent space learned by DC-gra2seq on DS3. (b) In each row, we morph sketch patterns by interpolating their corresponding latent codes.
Table 4: Sketch healing performance (%) on DS1. “Mask” denotes the probability for creating masks.
Mask Model $Rec\uparrow$ $Ret\uparrow$
@1 @10 @50
10% SketchHealer [1] 70.38 14.25 27.91 45.51
SketchHealer 2.0 [2] 90.18 19.45 41.51 59.94
SketchLattice [20] 54.18 1.09 5.48 14.00
SketchLattice+ 89.98 19.86 43.59 63.70
Lmser-pix2seq [14] 94.08 34.96 60.18 77.42
SP-gra2seq [7] 92.90 41.24 65.74 80.16
DC-gra2seq 94.45 74.59 88.83 94.31
@@
SketchHealer 2.0 [2] 79.05 4.66 14.54 28.44
SketchLattice [20] 32.03 0.08 2.05 5.96
SketchLattice+ 70.91 1.02 5.26 14.28
Lmser-pix2seq [14] 83.12 9.10 27.18 43.27
SP-gra2seq [7] 84.85 12.87 29.39 45.58
DC-gra2seq 86.50 30.30 51.38 65.84
Table 5: Sketch healing performance (%) on DS2.
Mask Model $Rec\uparrow$ $Ret\uparrow$
@1 @10 @50
10% SketchHealer [1] 70.56 35.22 55.86 66.24
SketchHealer 2.0 [2] 87.07 39.40 64.01 81.20
SketchLattice [20] 52.97 1.34 6.86 17.41
SketchLattice+ 90.52 46.98 70.21 83.12
Lmser-pix2seq [14] 92.20 53.99 73.76 83.61
SP-gra2seq [7] 91.24 50.42 73.35 85.18
DC-gra2seq 92.64 67.61 84.31 91.17
[…]
SketchHealer 2.0 [2] 75.08 10.05 24.51 39.57
SketchLattice [20] 31.73 0.88 3.64 9.25
SketchLattice+ 81.76 10.43 26.44 42.90
Lmser-pix2seq [14] 82.00 10.82 29.02 48.34
SP-gra2seq [7] 82.85 12.19 29.37 46.53
DC-gra2seq 84.62 29.92 49.26 62.44
Table 6: Sketch healing performance (%) on DS3.
Mask Model $Rec\uparrow$ $Ret\uparrow$
@1 @10 @50
10% SketchHealer [1] 60.91 15.10 37.89 53.10
SketchHealer 2.0 [2] 76.45 15.55 39.99 61.67
SketchLattice [20] 44.14 0.71 4.19 11.36
SketchLattice+ 78.28 22.56 44.89 62.11
Lmser-pix2seq [14] 83.75 35.99 52.31 69.85
SP-gra2seq [7] 83.38 40.20 63.40 77.52
DC-gra2seq 85.21 71.66 86.34 92.35
[…]
SketchHealer 2.0 [2] 60.74 3.75 11.66 22.36
SketchLattice [20] 23.06 0.41 2.41 5.96
SketchLattice+ 67.31 2.75 11.13 23.68
Lmser-pix2seq [14] 68.45 4.03 13.85 25.32
SP-gra2seq [7] 71.07 5.65 17.40 32.90
DC-gra2seq 72.84 27.64 47.34 60.61
@@ 4.3 Sketch healing
Table 4-6 present the sketch healing results on three datasets, respectively. Our DC-gra2seq yields significant improvements on $Ret$. When dealing with sketch patches corrupted by masks, the passing messages from the semantic-related patches by the values of cosine similarity could contain valuable missing information under the masks. Moreover, by taking account of PEs, DC-gra2seq might realize where these semantic-related patches are drawn from the contextual positions in a drawing order. By incorporating both visual patterns from the canvas and their sequential information from the embedded PEs, DC-gra2seq learns much more comprehensive and accurate sketch representations.
Figure 5: Qualitative comparisons on sketch healing.
4.4 Variant-drawing-protected DC-gra2seq
@@
Table 8: Controllable sketch synthesis (Mask = 0%) and sketch healing (Mask = 10%, 30%) performances of DC-gra2seq on DS1, by allowing PEs to participate in the construction of graph nodes, edges or not. “Node $\checkmark$ Edge $\times$” denotes the proposed DC-gra2seq.
Mask Node Edge $Rec$ $Ret$
@1 @10 @50
0% $\times$ $\times$ 94.51 80.90 95.09 98.19
$\checkmark$ $\times$ 96.26 96.84 99.48 99.88
$\times$ $\checkmark$ 95.69 79.22 93.92 97.77
$\checkmark$ $\checkmark$ 95.91 95.66 99.23 99.78
@@
$\checkmark$ $\times$ 94.45 74.59 88.83 94.31
$\times$ $\checkmark$ 93.48 50.64 75.10 87.46
$\checkmark$ $\checkmark$ 93.96 70.56 86.10 92.56
@@
$\checkmark$ $\times$ 86.50 30.30 51.38 65.84
$\times$ $\checkmark$ 84.23 16.28 36.80 55.01
$\checkmark$ $\checkmark$ 85.00 26.15 45.94 61.06
Table 9: Controllable sketch synthesis (Mask = 0%) and sketch healing (Mask = 10%, 30%) performances of DC-gra2seq on DS2, by allowing PEs to participate in the construction of graph nodes, edges or not.
Mask Node Edge $Rec$ $Ret$
@1 @10 @50
0% $\times$ $\times$ 85.84 41.00 71.80 87.31
$\checkmark$ $\times$ 95.70 92.75 98.08 99.42
$\times$ $\checkmark$ 92.41 67.49 88.30 95.30
$\checkmark$ $\checkmark$ 94.67 90.00 97.88 99.32
[…]
$\checkmark$ $\times$ 92.64 67.61 84.31 91.17
$\times$ $\checkmark$ 89.48 45.24 70.13 83.33
$\checkmark$ $\checkmark$ 91.55 67.10 84.08 91.02
[…]
$\checkmark$ $\times$ 84.62 29.92 49.26 62.44
$\times$ $\checkmark$ 80.40 18.62 37.78 53.70
$\checkmark$ $\checkmark$ 83.94 29.38 48.55 62.22
Table 10: Controllable sketch synthesis (Mask = 0%) and sketch healing (Mask = 10%, 30%) performances of DC-gra2seq on DS3, by allowing PEs to participate in the construction of graph nodes, edges or not.
Mask Node Edge $Rec$ $Ret$
@1 @10 @50
0% $\times$ $\times$ 83.93 57.50 83.52 93.45
$\checkmark$ $\times$ 90.41 96.27 99.47 99.81
$\times$ $\checkmark$ 88.48 76.25 92.64 97.24
$\checkmark$ $\checkmark$ 89.05 94.74 99.44 99.75
[…]
$\checkmark$ $\times$ 85.21 71.66 86.34 92.35
$\times$ $\checkmark$ 84.79 46.77 70.37 82.90
$\checkmark$ $\checkmark$ 85.18 69.96 86.67 92.43
[…]
$\checkmark$ $\times$ 72.84 27.64 47.34 60.61
$\times$ $\checkmark$ 71.99 14.56 31.54 47.96
$\checkmark$ $\checkmark$ 72.26 25.88 46.33 59.82
@@ 4.5 The impact of injecting drawing orders into nodes or edges
We further discuss the impact of injecting sketch drawing orders into graph nodes, edges or both via PEs. To make PEs access to edge construction, we allow the absolute PEs $\bm{P}$ to participate in the computation of the coefficients $\alpha_{t}$ directly, i.e., replacing Eq. (4) by $\alpha_{t}(i,j)=(\bm{v}_{ti}+\bm{P}(i))^{\top}(\bm{v}_{tj}+\bm{P}(j))$. A softmax activation function is applied over $\bm{A}_{t}$ in Eq. (5) before sent into Eq. (7) for message aggregation.
Table 8-10 present the experimental results on three datasets, respectively. When PEs attend the edge construction only (Node $\times$, Edge $\checkmark$), it outperforms the version which never uses PEs (Node $\times$, Edge $\times$) in most cases. The usage of drawing orders via PEs embedded graph edges is beneficial to improve sketch learning, but the performance gain is sensitive to various sketch drawings, e.g., it significantly outperforms “Node $\times$, Edge $\times$” on $Rec$ in DS2, but is defeated on top-1, top-10 $Ret$ in DS1. Similarly, our DC-gra2seq (Node $\checkmark$, Edge $\times$) outperforms “Node $\checkmark$, Edge $\checkmark$” in most cases. As a result, embedding drawing orders into graph nodes is a proper way to make better use of sequential information.
Table 11: Comparisons between variants of DC-gra2seq on DS1, by activating two types of PEs, or not. Mask = 0% and Mask = 10%, 30% indicate controllable sketch synthesis and sketch healing, respectively.
Mask Absolute PE Relative PE $Rec$ $Ret$
@1 @10 @50
0% $\times$ $\times$ 95.91 94.88 99.11 99.72
$\checkmark$ $\times$ 96.14 95.61 99.39 99.84
$\times$ $\checkmark$ 95.90 94.78 99.13 99.63
$\checkmark$ $\checkmark$ 96.26 96.84 99.48 99.88
[…]
$\checkmark$ $\times$ 94.17 72.45 88.38 93.92
$\times$ $\checkmark$ 93.61 59.17 78.50 88.34
$\checkmark$ $\checkmark$ 94.45 74.59 88.83 94.31
[…]
$\checkmark$ $\times$ 86.46 30.10 50.84 65.47
$\times$ $\checkmark$ 86.33 21.66 39.78 55.97
$\checkmark$ $\checkmark$ 86.50 30.30 51.38 65.84
Table 12: Comparisons between variants of DC-gra2seq on DS2, by activating two types of PEs, or not.
Mask Absolute PE Relative PE $Rec$ $Ret$
@1 @10 @50
0% $\times$ $\times$ 94.85 90.83 98.29 99.08
$\checkmark$ $\times$ 95.36 91.92 98.04 99.06
$\times$ $\checkmark$ 94.80 90.82 97.47 98.44
$\checkmark$ $\checkmark$ 95.70 92.75 98.08 99.42
[…]
$\checkmark$ $\times$ 92.45 66.11 84.01 90.57
$\times$ $\checkmark$ 91.89 63.92 81.44 90.04
$\checkmark$ $\checkmark$ 92.64 67.61 84.31 91.17
[…]
$\checkmark$ $\times$ 84.48 29.76 49.26 62.18
$\times$ $\checkmark$ 82.95 29.25 48.76 61.48
$\checkmark$ $\checkmark$ 84.62 29.92 49.26 62.44
Table 13: Comparisons between variants of DC-gra2seq on DS3, by activating two types of PEs, or not.
Mask Absolute PE Relative PE $Rec$ $Ret$
@1 @10 @50
0% $\times$ $\times$ 89.83 94.05 98.72 99.57
$\checkmark$ $\times$ 89.97 94.94 98.74 99.71
$\times$ $\checkmark$ 89.82 94.71 98.44 99.62
$\checkmark$ $\checkmark$ 90.41 96.27 99.47 99.81
[…]
$\checkmark$ $\times$ 94.09 58.36 76.07 85.51
$\times$ $\checkmark$ 83.88 49.40 69.86 81.77
$\checkmark$ $\checkmark$ 85.21 71.66 86.34 92.35
[…]
$\checkmark$ $\times$ 72.52 18.44 33.39 47.02
$\times$ $\checkmark$ 72.11 15.54 30.79 45.89
$\checkmark$ $\checkmark$ 72.84 27.64 47.34 60.61
4.6 Performance gained from two types of PEs
In this section, we verify the impacts of absolute and relative PEs introduced in DC-gra2seq. We deactivate two types of PEs by removing the matrices $\tilde{\bm{P}}$ or $\tilde{\bm{R}}$ in Eq. (10). The experimental results for three datasets are shown in Table 11-13.
With the equipments of both PEs, DC-gra2seq achieves the best performance in four variant models on controllable synthesis (Mask = 0%) and sketch healing (Mask = 10%, 30%). Moreover, it is worth noting that the attendance of absolute PE achieves significant improvements in both $Rec$ and $Ret$. The sequential information stored in drawing orders could be beneficial for learning graphic sketch representations, and a proper usage (e.g., injecting it into graph nodes by DC-gra2seq) could be more efficient to obtain the performance gain. Besides, activating relative PE jointly with absolute PE, rather than using relative PE solo, would contribute to better performances. The cooperation between two types of PEs guides DC-gra2seq focus on not only the pairwise sequential position differences between sketch patches, but also the information about when a patch is drawn in the drawing order. Moreover, our learnable relative PE encourages DC-gra2seq to flexibly utilize the contextual distances between patches, driving to performance improvements in both tasks.